OpenStudy (yacoub1993):

Solve the linear equation. 2(x + 6) + 6 = 3(x + 5) + 7 A. {6} B. {-4} C. {12} D. {9}

OpenStudy (yacoub1993):

i am solving and getting -18... how?

OpenStudy (yacoub1993):

@ranga help me please

OpenStudy (anonymous):

distribute to get 2x + 12 +16= 3x + 15 +7 simplify 2x + 28= 3x + 22 subtract 2x from each side and get 28=x+22 and then you finish with x= 6

OpenStudy (ranga):

^^^^ 2x + 12 + 6 (not 16)

OpenStudy (yacoub1993):

ohh i made a mistake in the middle solving

OpenStudy (anonymous):

rangs right idk what i was doing but its -4

OpenStudy (yacoub1993):

Find the product. (7x - 6)^2 A. 49x^2 + 36 B. 49x^2 - 84x + 36 C. 7x^2 - 84x + 36 D. 7x^2 + 36

OpenStudy (yacoub1993):

A. is correct right

OpenStudy (anonymous):

expand (7x - 6)^2 into (7x - 6) (7x - 6) then foil it to get the answer

OpenStudy (yacoub1993):

i didnt get ...then

OpenStudy (anonymous):

(7x - 6) (7x - 6) 49x^2 - 42x-42x + 36 49x^2 -84x +36

OpenStudy (yacoub1993):

from were did u get 84x

OpenStudy (anonymous):

-42x-42x gives you -84x

OpenStudy (anonymous):